6. A machine produces 9 inch latex gloves. A sample of 81 gloves is selected, and it is found that 37 are shorter than they should be. Find the 99% confidence interval on the proportion of all such gloves that are shorter than 9 inches
Answer by stanbon(75887)

A machine produces 9 inch latex gloves. A sample of 51 gloves is selected, and it is found that 46 are shorter than they should be.
Find the 98% confidence interval on the proportion of all such gloves that are shorter than 9 inches.
------------------------
sample proportion = 46/51 = 0.90
---
standard error = E = invNorm(0.985)*sqrt[0.9*0.1/(51)] = 0.04
----
98% CI: 0.9-0.04 < p < 0.9+0.04
98% CI: 0.86 < p < 0.94
===========================
Cheers,
Stan H.